Reports 1-1 of 1 Clear search Modify search
MIF (General)
keiko.kokeyama - 19:03 Thursday 19 December 2019 (12234) Print this report
SR3 jump, dither off, GR Y arm low trans

At 17:30 JST on Dec 19, SR3 suddenly jump in PIT for more than 2 urad and the GR TRANS dropped to 0.4. From 17:30-17:40, I tried to move SR3 back but stopped as it was still OBSERVATION mode. Maybe DA/Detchar people want to exculude this segment (but it shouldn't affect for IR or the sensitivity did not change). Since GR TRANS dropped (now only 0.1) SR3 dithering stopped too.

So it is still in OBSERVATION mode with SR3 still misaligned, SR3 dithering off.

Before the next lock acquisition, it should be put back. I gave TakahashiR-san and Kozakai-san (next operators) some instructions. I also write some notes here. Before the next lock acquisiton starts please:

  • Make sure VIS_SR3 guardian is ALIGNED. You can check on terminal by
    • guardctrl status VIS_SR3
  • Please put -26 to K1:VIS-SR3_IM_OLSET_P_OFFSET. From terminal, you cando
    • caput K1:VIS-SR3_IM_OLSET_P_OFFSET -26
  • Make sure K1:VIS-SR3_TM_OPLEV_PIT_DIAGMON goes to -26. You can check on terminal by
  • Make sure green Y arm transmission goes close to 1. Check by
    • caget K1:AOS-TMSY_GR_PDA1_OUTPUT
  • Then you can start LSC_LOCK guardian, or the initial alignment (if it seems necessary)
Images attached to this report
Comments to this report:
tomotada.akutsu - 19:31 Thursday 19 December 2019 (12235) Print this report

Altough GrY is zero in the DC-PD read, the gige camera at Y-end shows the transmitted green spot... I'm just surprised.

keiko.kokeyama - 21:45 Thursday 19 December 2019 (12238) Print this report

[SOLVED] TakahashiR-san put it back. 

ryutaro.takahashi - 21:47 Thursday 19 December 2019 (12239) Print this report

After the lock was lost, I put -26 to K1:VIS-SR3_IM_OLSET_P_OFFSET around pm8:30. The green Y arm transmission went to around 0.9. Then the LSC lock was recovered.

Search Help